Solve the porportion


16. 40/m = 32/16
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

16) 40/M=32/16 CROSS MULTIPLY
32M=40*16
32M=640
M=640/32
M=20 ANSWER.